if theta = 45, then whats the value of cosecant square theta?

Answer by sachi(548)

if theta = 45, then sin theta=sin 45=1/sqrt2
cosec theta=1/sin theta=sqrt2
the value of cosecant square theta=(sqrt2)^2=2